# Hotspot-Driven Peptide Design via Multi-Fragment Autoregressive Extension

### v1

@@ 5.3 Analysis
Effect of Hot Spots Comparing Tables  1 and  2, we observe that introducing hot spots as prior knowledge significantly boosts PepHAR’s performance, while providing little benefit to RFDiffusion and PepFlow. This highlights PepHAR’s versatility across different design tasks. We also investigate the effect of varying the number of hot spots, denoted as $K=1,2,3$. As shown in Tables and Figure  4, increasing the number of hot spots improves geometries and energies, regardless of whether the hotspots are estimated by density models or provided as ground truth; however, it negatively impacts novelty and diversity. This illustrates a trade-off between designing low-diversity but high-quality peptides (in comparison to the native) and high-diversity but varied peptides (Luo et al., 2022; Li et al., 2024a).
Ablation Study Table  3 presents our ablation study, which assesses the effectiveness of different components in PepHAR. ”PepHAR w/o Hot Spot” refers to the model where hot spots sampled from the density model are replaced with randomly positioned and typed residues. ”PepHAR w/o Von Mises” indicates the use of direct angle predictions instead of modeling angle distributions. We also remove the correction stage in ”PepHAR w/o Correction.” Our findings reveal that generated hot spots are crucial for Valid, RMSD, SSR, and BSR metrics, underscoring their importance for achieving valid geometries and interactions. Modeling angle distributions also contributes positively by accounting for the flexibility of dihedral angles. Lastly, the final correction stage plays a vital role in enhancing fragment assembly, leading to peptides with higher affinity and stability, which are essential for effective protein binding.
@@ 6 Conclusion
In this work, we presented PepHAR, a hot-spot based autoregressive generative model designed for efficient and precise peptide design targeting specific proteins. By addressing key challenges in peptide design—such as the unequal contribution of residues, the geometric constraints imposed by peptide bonds, and the need for practical benchmarking scenarios—PepHAR provides a comprehensive approach for generating peptides from scratch or assembling peptides around key hot spot residues. Our method leverages energy-based hot spot sampling, autoregressive fragment extension through dihedral angles, and an optimization process to ensure valid peptide assembly. Through extensive experiments on both peptide generation and scaffold-based design, we demonstrated the effectiveness of PepHAR in computational peptide design, highlighting its potential for advancing drug discovery and therapeutic development.

### v2

@@ 5.3 Analysis
Effect of Hot Spots Comparing Tables  1 and  2, we observe that introducing hot spots as prior knowledge significantly boosts PepHAR’s performance while providing little benefit to RFDiffusion and PepFlow. This highlights PepHAR’s versatility across different design tasks. We also investigate the effect of varying the number of hot spots, denoted as $K=1,2,3$. As shown in Tables and Figure  4, increasing the number of hot spots improves geometries and energies, regardless of whether the hotspots are estimated by density models or provided as ground truth; however, it negatively impacts novelty and diversity. This illustrates a trade-off between designing low-diversity but high-quality peptides (in comparison to the native) and high-diversity but varied peptides (Luo et al., 2022; Li et al., 2024a).
Ablation Study Table  3 presents our ablation study, which assesses the effectiveness of different components in PepHAR. ”PepHAR w/o Hot Spot” refers to the model where hot spots sampled from the density model are replaced with randomly positioned and typed residues. ”PepHAR w/o Von Mises” indicates using direct angle predictions instead of modeling angle distributions. We also removed the correction stage from ”PepHAR w/o Correction.” Our findings reveal that generated hot spots are crucial for Valid, RMSD, SSR, and BSR metrics, underscoring their importance for achieving valid geometries and interactions. Modeling angle distributions also contributes positively by accounting for the flexibility of dihedral angles. Lastly, the final correction stage is vital in enhancing fragment assembly, leading to peptides with higher affinity and stability, which are essential for effective protein binding.
@@ 6 Conclusion
In this work, we presented PepHAR, a hot-spot-based autoregressive generative model designed for efficient and precise peptide design targeting specific proteins. By addressing key challenges in peptide design—such as the unequal contribution of residues, the geometric constraints imposed by peptide bonds, and the need for practical benchmarking scenarios—PepHAR provides a comprehensive approach for generating peptides from scratch or assembling peptides around key hot spot residues. Our method leverages energy-based hot spot sampling, autoregressive fragment extension through dihedral angles, and an optimization process to ensure valid peptide assembly. Through extensive experiments on both peptide generation and scaffold-based design, we demonstrated the effectiveness of PepHAR in computational peptide design, highlighting its potential for advancing drug discovery and therapeutic development.
Acknowledgements
This work was supported by the National Key Plan for Scientific Research and Development of China (2023YFC3043300) and China’s Village Science and Technology City Key Technology funding.